OpenStudy (anonymous):

Which expression is a polynomial? A. 3/ square root x B. 2/x +3x C. square root x - 6 D. x - square root 6

OpenStudy (anonymous):

which one do u think??

jagr2713 (jagr2713):

Sorry it disconnected :D

OpenStudy (anonymous):

\[A. \frac{ 3 }{ \sqrt{x} }\]

OpenStudy (anonymous):

B. \[\frac{ 2 }{ x } +3x\]

OpenStudy (anonymous):

\[\sqrt{x} - 6\]

OpenStudy (anonymous):

D. \[x - \sqrt{6}\]

TheSmartOne (thesmartone):

Do you know what a polynomial is?

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

@jagr2713

jagr2713 (jagr2713):

oh no one helped you yet :?

OpenStudy (anonymous):

no I be stuck on this on for a wild

jagr2713 (jagr2713):

what do you think the answer is?

OpenStudy (anonymous):

I am thinking A

OpenStudy (anonymous):

this is the next question but i can't put anymore in the blue box its making it eeor Express the perimeter of the triangle as a polynomial.  A.–11x – 5 B. 11x – 4 C. 10x – 5 D. 11x + 4